A cell using Zn/Zn+2 (aq, 1M) and H2 / H+1 (aq, 1M) generates a cell potential of + 0.67 V under standard conditions.


A. Write the half-cell reaction at the anode.

B. Write the half-cell reaction at the cathode.

C. Write the electrochemical cell notation.

D. Calculate the standard reduction potential at the anode.



1
Expert's answer
2021-04-30T04:52:32-0400

(A): "Zn^0 (s) = Zn^{2+} (aq)+ 2e^-"


(C): "2H^+(aq) + 2e^- = H_2(g)"



"Zn (s) + 2H^+ (aq) = Zn^{2+} + H_2(g)"

"E_{cell} = E^0(2H^+\/H_2) - \u0415^0(Zn^0\/Zn^{2+}) = 0.763 V"